2 What is the present value of the following cash flow stream at a rate of 8.0%?







Years: 0123
 CFs: $750$2,450$3,175$4,400
Answer #1

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=750+2450/1.08+3175/1.08^2+4400/1.08^3

which is equal to

=$9233.43(Approx).
Answer #2

=

Present value=Cash flows*Present value of discounting factor(rate%,time period)

=750+2450/1.08+3175/1.08^2+4000/1.08^3

which is equal to

=$9233.43(Approx).
Answer #3

=750+2450/1.08+3175/1.08^2+4000/1.08^3